OpenStudy (anonymous):

The roots of y in the equation y2−2y=-9/2 are ?

OpenStudy (anonymous):

are you familiar with the quadratic equation, or completing the square?

OpenStudy (anonymous):

no i am not fimiliar

OpenStudy (anonymous):

Complete the square or use quadratic formula.

OpenStudy (anonymous):

i dont know what complete the square is.

OpenStudy (anonymous):

Are you learning about imaginary numbers?

OpenStudy (anonymous):

somewhat yes.

OpenStudy (anonymous):

ok is that a y2 or a \[y^{2}\]

OpenStudy (anonymous):

y^2

OpenStudy (anonymous):

ok lets use the Quadratic formula

OpenStudy (anonymous):

a=1 b=2 c=-9/2

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4(a)(c)} }{2 (a) }\]

OpenStudy (anonymous):

shouldnt b=-2?

OpenStudy (anonymous):

yes sorry i messed up

OpenStudy (anonymous):

a=1 b=-2 c=-9/2

OpenStudy (anonymous):

\[\frac{ 2\pm \sqrt{2^{2}-4(1)(\frac{ -9 }{ 2 })} }{ 2(1) }\]

OpenStudy (anonymous):

okay this is where i start to get confused

OpenStudy (anonymous):

ok lets work throughout it by sections

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lets start here \[\sqrt{2^{2}-4(1)(-\frac{ 9 }{ 2 })}\]

OpenStudy (anonymous):

\[\sqrt{4+18}\]

OpenStudy (anonymous):

\[\sqrt{22}\]

OpenStudy (anonymous):

\[\sqrt{22}\]

OpenStudy (anonymous):

\[\frac{ 2 \pm \sqrt{22} }{ 2 }\]

OpenStudy (anonymous):

okay im not sure what this equals

OpenStudy (anonymous):

I'm sorry i messed up again the -9/2 should have turned positive so it would have been \[\sqrt{4-18}\]

OpenStudy (anonymous):

\[\sqrt{-14}\]

OpenStudy (anonymous):

\[\frac{ 2\pm \sqrt{-14} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 2\pm \iota \sqrt{14} }{ 2 }\]

OpenStudy (anonymous):

that is as low as it will go

OpenStudy (anonymous):

okay thanks so much!!!

OpenStudy (anonymous):

no problem!